# TrustChain: A Blockchain Framework for Auditing and Verifying Aggregators in Decentralized Federated Learning

### v1

@@ Abstract
The server-less nature of Decentralized Federated Learning (DFL) requires allocating the aggregation role to specific participants in each federated round. Current DFL architectures ensure the trustworthiness of the aggregator node upon selection. However, most of these studies overlook the possibility that the aggregating node may turn rogue and act maliciously after being nominated. To address this problem, this paper proposes a DFL structure, called TrustChain, that scores the aggregators before selection based on their past behavior and additionally audits them after the aggregation. To do this, the statistical independence between the client updates and the aggregated model is continuously monitored using the Hilbert-Schmidt Independence Criterion (HSIC). The proposed method relies on several principles, including blockchain, anomaly detection, and concept drift analysis. The designed structure is evaluated on several federated datasets and attack scenarios with different numbers of Byzantine nodes.
Index Terms:
Federated learning, blockchain, Byzantine attacks, data poisoning, model poisoning, security.
@@ I Introduction
To address this gap, we propose a novel approach, called TrustChain, that leverages blockchain technology to ensure the continuous trustworthiness of the aggregator node throughout the federated learning process. This method consists of two main components:
@@ II-A Decentralized Federated Learning
DFL primarily addresses the issue of the single point of failure in FL, by removing the dependency on a reliable server to carry out the aggregation and enhance communication efficiency [15, 16]. On the client side, the process of training the local models and preparing the updates for the global aggregator is similar to that of FL. However, due to the lack of a central server in DFL, exchanging model parameters and model aggregation are performed using peer-to-peer communications or blockchain technology [10]. Integration with blockchain provides additional benefits such as traceability and immutability. Despite DFL’s security and efficiency characteristics, the design is not faultless. For example, adding blockchain into FL may leave the system vulnerable to blockchain-related security threats. More importantly, the trustworthiness of participants is more critical in DFL, as the aggregating nodes are selected among them in each round.
@@ II-B Threats
Similar to federated learning, threats to DFL generally fall under privacy and performance-related attacks [10]. Privacy attacks primarily aim at stealing sensitive data during inference based on the communicated gradients. An example of privacy attacks in DFL is presented by [13]. On the other hand, poisoning attacks target the global model by injecting poisonous updates during the training. The attacker can use data and model poisoning to deteriorate the overall performance of the global model. For instance, [6] performs data poisoning to disrupt the global performance. This study assumes 30 percent of the nodes are malicious, and considers that this ratio remains unchanged in the experiments. The same objective can be achieved using Gradient Manipulation (GM), in which Byzantine data holders generate forged gradients sampled from a Gaussian distribution [17]. The same malicious objective is satisfied in [18] by filliping the class labels randomly on the training data. Moreover, DFL can be exposed to more subtle attacks such as backdoors as presented in [13, 19].
@@ II-C Defenses
Biscotti was among the first efforts to secure DFL against privacy attacks through blockchain integration [6]. The method integrates Differential Privacy (DP) with secure aggregation as a defense against poisoning and inference attacks. Another study proposes LearningChain, a DFL structure that also uses DP and an aggregation mechanism that works based on decentralized Stochastic Gradient Descent (SGD) to eliminate Byzantine attacks [17]. One of the first efforts to demonstrate the use of smart contracts in DFL was made by [8]. This method uses smart contracts to keep track of client states and store the global model. Proof of correctness can also be used by registered users to audit the generated updates [22]. [9] uses a blockchain-based committee consensus mechanism to enhance the privacy and efficiency of DFL. The sharding approach commonly used in blockchain is used by [23] to secure the aggregation process. [7] proposed VFChain, a structure that audits trainers and generates updates using a trusted committee. The committee members are selected arbitrarily, and they use a verifying contract to verify aggregated models. Another approach proposed by [13] integrates several methods such as DP, anomaly detection, and gradient pruning to safeguard DFL against poisoning attacks. [24] tackles the issue of malicious aggregators in DFL by using zero-knowledge proofs and blockchain to enable secure and verifiable aggregation.
@@ III Problem Formulation
In this scenario, the two main attack surfaces are the client updates and the aggregated parameters. Firstly, $\theta_{i}$ uploaded to $B$ may be poisoned to satisfy a malicious goal. These threats are common with the centralized architecture of federated learning and are commonly addressed using defense mechanisms such as robust aggregation and anomaly detection. Secondly, even if the updates are filtered and the malicious updates are dropped, the node aggregating the results $c^{*}$ can still compromise the entire network by updating the global state into a malicious set of parameters.
@@ IV Threat Models
While the spectrum of poisoning attacks is vast, this paper focuses on model poisoning attacks. We consider four poisoning attacks, namely label flipping, sign flipping, GM, and Training Objective Manipulation (TOM). When these attacks are initiated from the client side, a robust aggregator attempts to mitigate their effects. If a malicious node is selected as an aggregator, it can bypass the robust aggregator by directly feeding the poisoned parameters to the DFL. These attacks are formally explained as follows.
@@
The labels of the training data are flipped to obtain corrupted parameters [25]. Let $\mathcal{D}_{i}={(x_{i},y_{i})}$ be the dataset of node $c_{i}$, where $y_{i}$ is the true label. The attacker generates a poisoned dataset $\tilde{\mathcal{D}}_{i}$ by randomly flipping labels:
@@
~$\tilde{\mathcal{L}}(\bar{\theta}^{t})=\mathcal{L}(\bar{\theta}^{t})+\zeta\cdot\mathcal{P}(\bar{\theta}^{t}),$ (5)
@@ V Blockchain-Enabled Trustworthy Aggregation
Given that the goal of this study is to ensure the trustworthiness of the aggregator, the proposed TrustChain assumes the employed robust aggregation algorithm is reliable. Consequently, this research focuses on securing the aggregation phase by performing PSE and PAA in DFL. The process begins with the PSE algorithm evaluating the past behavior of participants using concept drift analysis to determine a set of trustworthy candidates. Then, the aggregated model is further evaluated by PAA based on the statistical independence of $\bar{\theta}^{t}$ with that of the previous block and $\theta_{i}$ received in the current round.
V-A Pre-Selection Evaluation
@@ 8             Calculate $\operatorname{CosSim}(\theta_{i}^{j-1},\bar{\theta}^{j})$
where $1\leq i\leq r$. Note that we estimate the similarity for all nodes that participated in round $t-1$ regardless of the fact that not all $\theta_{i}^{t-1}$ may have been used to estimate $\bar{\theta^{t}}$. This approach results in a lower score for $c_{i}$ whose $\theta_{i}$ are disregarded or not used during the aggregation. To analyze the node behavior in a longer time span, a window of $q$ blocks is used to reveal the overall drift $c_{i}$ creates in the path of $\bar{\theta}$. Correspondingly, a score is calculated for each node at each time step.
@@
V-B Post-Aggregation Auditing
@@
where ||$\cdot$|| ⋅ | returns the cardinality. In addition, $H=\mathbf{I}-\frac{1}{|\theta|}\mathbf{11}^{\top}$ is the centring matrix, and $F_{A}=f_{A}(a_{i},a_{j})$ and $F_{B}=f_{B}(b_{i},b_{j})$ are kernel matrices obtained from $a_{i}\in A$ and $b_{i}\in B$.
@@ 14       Trigger $SC_{1}$ to select a new aggregator
V-C Blockchain Components
@@ VI-A Experimental Setting
Datasets
Experiments are evaluated using commonly used benchmark datasets in this domain, such as MNIST [29], Fashion-MNIST [30], and CIFAR10 [31] datasets. To create federated datasets, datasets are shuffled at first and then sub-datasets are created by random sampling without replacement. The resulting data is highly imbalanced, and the created subsets have an equal number of samples.
Local models
The architecture of local models differs for each dataset. For experiments on MNIST and Fashion-MNIST, a Multi-Layer Perceptron (MLP) with two hidden layers is used. For CIFAR10, on the other hand, we use a Convolutional Neural Network (CNN) with six convolutional layers followed by two dense layers. Convolution layers are paired with batch normalization, and after every two convolutional layers, max pooling and dropout (ratio=$25\%$) are used. In both networks, hidden layers use ReLU, and the final layer is activated using Softmax. Model parameters are optimized using stochastic gradient descent. The batch size is set to 64 for MLP and 32 for CNN.
DFL parameters
@@ VI-B Results Analysis
Figure 2: Classification accuracy of DFL under three different conditions: (a–c) aggregator is benign, (d–f) uses TrustChain and aggregator is not trusted, (g–i) TrustChain is not used and aggregator is not trusted. For each scenario, experiments are separated based on the utilized datasets.
Overall performance
@@
In order to determine the effectiveness of TrustChain, we first compare the classification accuracy of DFL under four different attacks. In these experiments, malicious users attempt to inject poisonous updates into the DFL network. When DFL is not paired with TrustChain, aggregator nodes are selected randomly. Since the DFL is equipped with a robust aggregator, the effect of these attacks is expected to be partially mitigated when not initiated by the aggregator. Here, we select TrimmedMean [32] for the task of robust aggregation. Nevertheless, when these nodes are selected as the aggregator, they will have the opportunity to bypass the robust aggregator. Fig. 2 compares the classification accuracy of TrimmedMean under three conditions when poisoning attacks are launched with different ratios of Byzantine nodes, ranging from five to twenty percent of the nodes’ population. To identify a baseline, we first investigate a scenario where the aggregator is benign and TrustChain is not used in the DFL, as shown in panels (a–c) of Fig. 2. This shows the performance of TrimmedMean in attack scenarios, without a malicious aggregator becoming involved. As expected, the accuracy deteriorates as the ratio of Byzantine nodes increases. Next, we assume the aggregator is not trusted, that is a malicious node can be chosen to carry out the aggregation. Fig. 2(d–f) shows the evaluation results of TrustChain under this condition. The accuracy obtained in these panels is slightly lower than that of Fig. 2(a–c). This indicates that TrustChain can robustly mitigate malicious aggregators, albeit not completely. To reveal the effectiveness of TrustChain in this scenario, we remove it from the experiments in Fig. 2(g–i) to demonstrate the potency of malicious aggregators when they are not audited. As illustrated, a noticeable performance drop has taken place as the bypassed poisoned update rapidly propagates through the DFL network. This effect becomes more pronounced as the ratio of Byzantine nodes increases because malicious aggregators will have a higher chance of being selected in each round.
Sensitivity to attacks

### v2

@@ Abstract
The serverless nature of Decentralized Federated Learning (DFL) requires allocating the aggregation role to specific participants in each federated round. Current DFL architectures ensure the trustworthiness of the aggregator node upon selection. However, most of these studies overlook the possibility that the aggregating node may turn rogue and act maliciously after being nominated. To address this problem, this paper proposes a DFL structure, called TrustChain, that scores the aggregators before selection based on their past behavior and additionally audits them after the aggregation. To do this, the statistical independence between the client updates and the aggregated model is continuously monitored using the Hilbert-Schmidt Independence Criterion (HSIC). The proposed method relies on several principles, including blockchain, anomaly detection, and concept drift analysis. The designed structure is evaluated on several federated datasets and attack scenarios with different numbers of Byzantine nodes.
@@ I Introduction
To address this gap, we propose a novel approach, called TrustChain, that leverages blockchain technology to ensure the continuous trustworthiness of the aggregator node throughout the DFL process. This method consists of two main components:
@@ II-A Decentralized Federated Learning
DFL primarily addresses the issue of the single point of failure in FL, by removing the dependency on a reliable server to carry out the aggregation and enhance communication efficiency [15, 16]. On the client side, the process of training the local models and preparing the updates for the global aggregator is similar to that of FL. However, due to the lack of a central server in DFL, exchanging model parameters and model aggregation are performed using peer-to-peer communications or blockchain technology [10]. Integration with blockchain provides additional benefits such as traceability and immutability [17, 18]. Despite DFL’s security and efficiency characteristics, the design is not faultless. For example, adding blockchain into FL may leave the system vulnerable to blockchain-related security threats. More importantly, the trustworthiness of participants is more critical in DFL, as the aggregating nodes are selected among them in each round.
@@ II-B Threats
Similar to FL, threats to DFL generally fall under privacy and performance-related attacks [10]. Privacy attacks primarily aim at stealing sensitive data during inference based on the communicated gradients. An example of privacy attacks in DFL is presented by [13]. On the other hand, poisoning attacks target the global model by injecting poisonous updates during the training. The attacker can use data and model poisoning to deteriorate the overall performance of the global model. For instance, [6] performs data poisoning to disrupt the global performance. This study assumes 30 percent of the nodes are malicious, and considers that this ratio remains unchanged in the experiments. The same objective can be achieved using Gradient Manipulation (GM), in which Byzantine data holders generate forged gradients sampled from a Gaussian distribution [19]. The same malicious objective is satisfied in [20] by filliping the class labels randomly on the training data. Moreover, DFL can be exposed to more subtle attacks such as backdoors as presented in [13, 21].
@@ II-C Defenses
Biscotti was among the first efforts to secure DFL against privacy attacks through blockchain integration [6]. The method integrates Differential Privacy (DP) with secure aggregation as a defense against poisoning and inference attacks. Another study proposes LearningChain, a DFL structure that also uses DP and an aggregation mechanism that works based on decentralized Stochastic Gradient Descent (SGD) to eliminate Byzantine attacks [19]. One of the first efforts to demonstrate the use of smart contracts in DFL was made by [8]. This method uses smart contracts to keep track of client states and store the global model. Proof of correctness can also be used by registered users to audit the generated updates [24]. The sharding approach commonly used in blockchain is used by [25] to secure the aggregation process. [7] proposed VFChain, a structure that audits trainers and generates updates using a trusted committee. The committee members are selected arbitrarily, and they use a verifying contract to verify aggregated models. Another approach proposed by [13] integrates several methods such as DP, anomaly detection, and gradient pruning to safeguard DFL against poisoning attacks. Zero-Knowledge (ZK) proofs is used in [26] so that each round’s aggregation is verifiably faithful. To do so, the aggregator generates a proof of correct model aggregation, which blockchain miners validate without exposing client models, ensuring both integrity and privacy. VerifBFL extends this concept with ZK Succinct Non-interactive Arguments of Knowledge (ZK-SNARKs) and incrementally verifiable computation, allowing proofs to be generated for both local training and aggregation steps, which are then verified on-chain in under a second, offering trustless correctness but at the cost of cryptographic overhead [27]. Secure Verifiable Aggregation frameworks combine public verifiable secret sharing with masking and encryption to hide individual updates while still enabling blockchain participants to verify the correctness of the aggregated model, providing dropout resilience in addition to integrity guarantees [28]. VDFChain employs verifiable delay functions and polynomial commitments to make aggregated updates verifiable while deterring manipulation [29]. Committee-based consensus designs, such as BFLC, assign aggregation to a selected committee whose results are recorded on-chain, reducing reliance on a single aggregator and mitigating certain attacks [9]. Verifiable off-chain computation frameworks also use ZK-proof systems to execute aggregation off-chain while producing proofs that are verified on-chain, reducing latency while maintaining integrity [30]. BFL-SA focuses on enhancing secure aggregation in blockchain-based FL, protecting the integrity and confidentiality of aggregated updates with lightweight cryptographic protocols [31]. To secure DFL in IoT, [32] designs a hybrid node selection mechanism that combines reputation scores with reinforcement learning, and a multi-level aggregation scheme to improve reliability and handle asynchronicity. While this indirectly enhances aggregator trust by prioritizing capable nodes, their approach does not verify aggregator behavior after aggregation.
II-D Knowledge Gap
While recent studies have explored various blockchain-based mechanisms to enhance the security of DFL, most existing solutions secure the aggregator either through cryptographic verifiability or trusted node selection. Cryptographic proof-based methods ensure that aggregation is performed correctly, but they incur significant computational and communication overhead, require specialized proof systems, and may be challenging to scale in large, resource-constrained deployments. Trusted selection approaches indirectly improve aggregator trustworthiness by reducing the likelihood of selecting a malicious node. However, these methods do not verify aggregator behavior after aggregation and thus cannot detect malicious actions from otherwise trusted aggregators. Furthermore, several works focus on complementary issues, such as privacy preservation, dropout resilience, and asynchronicity handling, without providing continuous, fine-grained verification of aggregation integrity throughout training rounds.
This reveals a clear gap: there is a lack of lightweight, topology-agnostic mechanisms that can continuously monitor and verify the trustworthiness of the aggregator’s behavior during the learning process, without relying on heavy cryptography or fixed committee structures. TrustChain addresses this gap through a behavior-based trust framework that is adaptable to various FL settings and complementary to proof-based methods. Its PSE filters aggregators before selection using historical performance patterns, while its PAA statistically analyzes aggregated models for anomalies after aggregation. This combination provides continuous, low-overhead security that can operate independently or be integrated with existing blockchain-based FL frameworks for layered protection.
@@ III Problem Formulation
In this scenario, the two main attack surfaces are the client updates and the aggregated parameters. Firstly, $\theta_{i}$ uploaded to $B$ may be poisoned to satisfy a malicious goal. These threats are common with the centralized architecture of FL and are commonly addressed using defense mechanisms such as robust aggregation and anomaly detection. Secondly, even if the updates are filtered and the malicious updates are dropped, the node aggregating the results $c^{*}$ can still compromise the entire network by updating the global state into a malicious set of parameters.
@@ IV Threat Models
While the spectrum of poisoning attacks is vast, this paper focuses on model poisoning attacks. We consider four poisoning attacks, namely label flipping, sign flipping, GM, and Training Objective Manipulation (TOM). These attacks were selected because they represent both simple-yet-effective poisoning strategies (e.g., label/sign flipping) and more sophisticated attacks (e.g., GM, TOM) that can evade conventional defenses. For each attack, we vary the proportion of Byzantine nodes in the population (1–20%) to study performance under different threat intensities. Attacks are injected in two distinct ways: (i) Client-side attackers, where malicious nodes contribute poisoned local updates during training, and (ii) Aggregator-side attackers, where the selected aggregator is malicious and directly uploads a poisoned global model, bypassing the robust aggregation stage. The latter case is critical for evaluating TrustChain, as it explicitly targets our main security objective: detecting and mitigating rogue aggregators.
When these attacks are initiated from the client side, a robust aggregator attempts to mitigate their effects. If a malicious node is selected as an aggregator, it can bypass the robust aggregator by directly feeding the poisoned parameters to the DFL. These attacks are formally explained as follows.
@@
The labels of the training data are flipped to obtain corrupted parameters [33, 34]. Let $\mathcal{D}_{i}={(x_{i},y_{i})}$ be the dataset of node $c_{i}$, where $y_{i}$ is the true label. The attacker generates a poisoned dataset $\tilde{\mathcal{D}}_{i}$ by randomly flipping labels:
@@
$\tilde{\mathcal{L}}(\bar{\theta}^{t})=\mathcal{L}(\bar{\theta}^{t})+\zeta\cdot\mathcal{P}(\bar{\theta}^{t}),$ (5)
@@ V Blockchain-Enabled Trustworthy Aggregation
If DFL operates without an aggregator evaluation mechanism, aggregator nodes may be chosen using methods that do not account for their trustworthiness, which increases the risk of selecting a malicious node. In the proposed method, aggregator security is ensured through PSE and PAA mechanisms. PSE operates before selection and scores miners over a sliding window of $q$ previous rounds based on their historical alignment with the global model’s trajectory using concept drift analysis. This allows the system to prioritize the most trustworthy candidate and significantly reduce the likelihood of a rogue node being selected. PAA operates after aggregation and statistically examines the aggregated model to identify anomalies or signs of poisoning. If PSE is missing, the selection process lacks behavior-based filtering, increasing the chance that an untrustworthy aggregator is chosen. If PAA is missing, there is no safeguard to detect an aggregator that behaves maliciously after being selected, which allows a compromised node to poison the model without detection. Together, PSE and PAA form a layered defense where PSE filters aggregators before selection and PAA verifies their behavior afterward.
Given that the goal of this study is to ensure the trustworthiness of the aggregator, the proposed TrustChain assumes the employed robust aggregation algorithm is reliable. Consequently, this research focuses on securing the aggregation phase by performing PSE and PAA in DFL. The process begins with the PSE algorithm evaluating the past behavior of participants using concept drift analysis to determine a set of trustworthy candidates. Then, the aggregated model is further evaluated by PAA based on the statistical independence of $\bar{\theta}^{t}$ with that of the previous block and $\theta_{i}$ received in the current round. The following subsections describe each component in detail, with the final subsection dedicated to the complexity analysis.
V-A Robust Aggregation
In our experiments, we use the TrimmedMean algorithm as the robust aggregator [37]. Robust aggregation aims to mitigate the influence of poisoned or extreme client updates by limiting the effect of outliers in each model parameter coordinate. Let $\{\theta_{i}\}_{i=1}^{r}$ denote the set of full model parameter vectors received from $r$ clients in a round, where each $\theta_{i}$ is trained locally on $m$ samples. For each parameter coordinate $j$, TrimmedMean collects scalar values $\theta_{i}^{j}$ across all clients, sorts them, removes a fixed fraction of the largest and smallest values, and averages the remaining ones. This coordinate-wise trimming is applied independently for all dimensions of the model parameter vector, producing the aggregated model $\bar{\theta}$.
The choice of robust aggregator is not part of TrustChain’s novelty. TrimmedMean is used here solely as a representative baseline due to its simplicity and established Byzantine robustness, ensuring a fair and reproducible evaluation of TrustChain. However, more advanced robust aggregators can be substituted without altering the operation of PSE or PAA.
V-B Pre-Selection Evaluation
@@ 8      Calculate $\operatorname{CosSim}(\theta_{i}^{j-1},\bar{\theta}^{j})$
where $1\leq i\leq r$. Note that we estimate the similarity for all nodes that participated in round t1t-1 regardless of the fact that not all $\theta_{i}^{t-1}$ may have been used to estimate $\bar{\theta^{t}}$. This approach results in a lower score for $c_{i}$ whose $\theta_{i}$ are disregarded or not used during the aggregation. To analyze the node behavior in a longer time span, a window of $q$ blocks is used to reveal the overall drift $c_{i}$ creates in the path of $\bar{\theta}$. Correspondingly, a score is calculated for each node at each time step.
@@
PSE continuously monitors concept drift to ensure a candidate aggregator’s updates remain consistent with the evolving concept represented by the global model trajectory. Here, the concept refers to the statistical relationship between successive aggregated models and the local updates that contributed to them. For each node $c_{i}$, we track its historical behavior over the last $q$ rounds by measuring the cosine similarity between its submitted parameters $\theta_{i}^{t-1}$ and the corresponding aggregated model $\bar{\theta}^{t}$. This similarity captures both direction and magnitude alignment with the global model’s learning path. A sustained drop in similarity, especially when recent rounds are weighted more heavily via the decay factor $\alpha^{t-j}$, indicates a drift from the consensus learning trajectory, which may be due to malicious behavior (e.g., poisoning) or degraded participation quality. The aggregated score $S_{i}^{t}$ across the sliding window thus functions as a drift indicator so that high scores signal stable, aligned contributions, while low scores reveal significant concept drift. PSE uses this drift score to exclude candidates whose historical updates suggest a high likelihood of deviating from the global model’s objective, thereby reducing the risk of appointing a malicious or unreliable aggregator.
V-C Post-Aggregation Auditing
@@
where $|\cdot|$ returns the cardinality. In addition, $H=\mathbf{I}-\frac{1}{|\theta|}\mathbf{11}^{\top}$ is the centring matrix, and $F_{A}=f_{A}(a_{i},a_{j})$ and $F_{B}=f_{B}(b_{i},b_{j})$ are kernel matrices obtained from $a_{i}\in A$ and $b_{i}\in B$.
@@ 14   Trigger $SC_{1}$ to select a new aggregator
PSE and PAA are seperated intentionally. PSE acts early to filter out clearly harmful updates before aggregation, reducing the computational burden and limiting the influence of severe outliers, while PAA operates afterward on the aggregated model to catch subtler anomalies that cannot be detected beforehand. By separating PSE and PAA, we provided both efficiency, and a second line of defense of additional protection, which enhances its robustness.
V-D Blockchain Components
@@ V-D2 Smart contracts
V-E Time Complexity
Let $d$ denote the model parameters’ dimensionality, $k$ the number of candidate miners, $r$ the number of client updates in the current round, and $q$ the window size for historical statistics.
In PSE, the trust score of each candidate miner is computed as a decayed sum of cosine similarities over the last $q$ rounds. Computing a cosine similarity between two $d$-dimensional vectors requires a dot product and two norm calculations, each taking $O(d)$ operations, resulting in $O(d)$ per similarity. Since each of the $k$ candidates is evaluated over $q$ past rounds, PSE performs $kq$ such computations per round. Therefore, the per-round time complexity of PSE is $O(kqd)$.
PAA begins by computing the elementwise median of $r$ client updates, each of dimension $d$. This can be achieved in $O(rd)$ using linear-time selection for each coordinate. It then evaluates the HSIC between the median update and the proposed aggregated model. For a generic kernel, HSIC computation involves constructing $d\times d$ Gram matrices and centering them, leading to $O(d^{2})$ complexity. If a linear kernel is used, this reduces to $O(d)$. Finally, the threshold estimation step uses only $q$ stored HSIC values and requires $O(q)$, which is negligible compared to the other terms. Thus, the per-round complexity of PAA is $O(rd+d^{2}+q)$.
Combining both stages, the per-round time complexity of TrustChain is $O(kqd)+O(rd+d^{2}+q)=O(kqd+rd+d^{2})$. For high-dimensional models where $d\gg r$, the $O(d^{2})$ HSIC term dominates. If a linear kernel or approximation is applied, the overall complexity reduces to $O(kqd+rd)$, scaling linearly with the model size and number of candidates or participants.
@@ VI-A Experimental Setting
VI-A1 Datasets
Experiments are evaluated using commonly used benchmark datasets from different domains, such as MNIST [39], Fashion-MNIST [40], CIFAR10 [41], Shakespeare [1], and Credit Card [42] datasets. To create federated datasets, datasets are shuffled at first and then sub-datasets are created by random sampling without replacement. The resulting data is highly imbalanced, and the created subsets have an equal number of samples.
VI-A2 Local models
[…]
VI-A3 DFL parameters
@@ VI-B Results Analysis
Figure 2: Classification accuracy of DFL for different datasets. On the left side, aggregator is benign. Plots in the middle column use TrustChain and aggregator is not trusted. On the right side, DFL do not use TrustChain and aggregator is not trusted.
VI-B1 Overall performance
@@
In order to determine the effectiveness of TrustChain, we first compare the classification accuracy of DFL under four different attacks. In these experiments, malicious users attempt to inject poisonous updates into the DFL network. When DFL is not paired with TrustChain, aggregator nodes are selected randomly. Since the DFL is equipped with a robust aggregator, the effect of these attacks is expected to be partially mitigated when not initiated by the aggregator. Here, we select TrimmedMean [37] for the task of robust aggregation. Nevertheless, when these nodes are selected as the aggregator, they will have the opportunity to bypass the robust aggregator. Fig. 2 compares the classification accuracy of TrimmedMean under three conditions when poisoning attacks are launched with different ratios of Byzantine nodes, ranging from one to twenty percent of the nodes’ population. To identify a baseline, we first investigate a scenario where the aggregator is benign and TrustChain is not used in the DFL, as shown in panels on the left side of Fig. 2. This shows the performance of TrimmedMean in attack scenarios, without a malicious aggregator becoming involved. As expected, the accuracy deteriorates as the ratio of Byzantine nodes increases. Next, we assume the aggregator is not trusted, that is a malicious node can be chosen to carry out the aggregation. The middle panels in Fig. 2 shows the evaluation results of TrustChain under this condition. The accuracy obtained in these panels is slightly lower than that of the benign aggregator in Fig. 2. This indicates that TrustChain can robustly mitigate malicious aggregators, albeit not completely. To reveal the effectiveness of TrustChain in this scenario, we remove it from the experiments, shown in plots on the right side of Fig. 2, to demonstrate the potency of malicious aggregators when they are not audited. As illustrated, a noticeable performance drop has taken place as the bypassed poisoned update rapidly propagates through the DFL network. This effect becomes more pronounced as the ratio of Byzantine nodes increases because malicious aggregators will have a higher chance of being selected in each round.
VI-B2 Sensitivity to attacks